# Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection

### v1

@@ 4.2. RQ2: How do different components affect LLM4FL’s fault localization accuracy?
Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parenthesis show the percentage changes compared to LLM4FL with all the components.
       Project (# faults)        Techniques        Top-1        Top-3        Top-5        Top-10
       Cli (39)        LLM4FL        16        21        23        24
       LLM4FLw/o CodeNavigation        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
       LLM4FLw/o Divide&Conquer        11 (-31.25%)        17 (-19.05%)        17 (-26.09%)        19 (-20.83%)
       LLM4FLw/o Reflection        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
       LLM4FLw/o CoT        16 (0.0%)        21 (0.0%)        23 (0.0%)        24 (0.0%)
       Closure (174)        LLM4FL        52        77        102        118
       LLM4FLw/o CodeNavigation        32 (-38.46%)        45 (-41.56%)        53 (-48.04%)        53 (-55.08%)
       LLM4FLw/o Divide&Conquer        27 (-48.08%)        40 (-48.05%)        49 (-51.96%)        50 (-57.63%)
       LLM4FLw/o Reflection        50 (-3.85%)        74 (-3.9%)        92 (-9.8%)        109 (-7.63%)
       LLM4FLw/o CoT        52 (0.0%)        77 (0.0%)        102 (0.0%)        118 (0.0%)
       Codec (18)        LLM4FL        9        13        13        13
       LLM4FLw/o CodeNavigation        8 (-11.11%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
       LLM4FLw/o Divide&Conquer        8 (-11.11%)        11 (-15.38%)        12 (-7.69%)        13 (0.0%)
       LLM4FLw/o Reflection        9 (0.0%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
       LLM4FLw/o CoT        9 (0.0%)        13 (0.0%)        13 (0.0%)        13 (0.0%)
       Compress (47)        LLM4FL        23        32        34        34
       LLM4FLw/o CodeNavigation        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
       LLM4FLw/o Divide&Conquer        23 (0.0%)        28 (-12.5%)        30 (-11.76%)        31 (-8.82%)
       LLM4FLw/o Reflection        23 (0.0%)        32 (0.0%)        35 (2.94%)        35 (2.94%)
       LLM4FLw/o CoT        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
       Csv (16)        LLM4FL        8        10        10        10
       LLM4FLw/o CodeNavigation        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
       LLM4FLw/o Divide&Conquer        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
       LLM4FLw/o Reflection        7 (-12.5%)        10 (0.0%)        10 (0.0%)        10 (0.0%)
       LLM4FLw/o CoT        8 (0.0%)        9 (-10.0%)        10 (0.0%)        10 (0.0%)
       Gson (16)        LLM4FL        11        14        14        14
       LLM4FLw/o CodeNavigation        9 (-18.18%)        15 (7.14%)        15 (7.14%)        15 (7.14%)
       LLM4FLw/o Divide&Conquer        10 (-9.09%)        14 (0.0%)        15 (7.14%)        15 (7.14%)
       LLM4FLw/o Reflection        11 (0.0%)        12 (-14.29%)        13 (-7.14%)        14 (0.0%)
       LLM4FLw/o CoT        11 (0.0%)        14 (0.0%)        14 (0.0%)        14 (0.0%)
       JacksonCore (26)        LLM4FL        12        13        15        15
       LLM4FLw/o CodeNavigation        9 (-25.0%)        13 (0.0%)        14 (-6.67%)        14 (-6.67%)
       LLM4FLw/o Divide&Conquer        9 (-25.0%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
       LLM4FLw/o Reflection        10 (-16.67%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
       LLM4FLw/o CoT        12 (0.0%)        13 (0.0%)        15 (0.0%)        15 (0.0%)
       JacksonXml (6)        LLM4FL        4        4        4        4
       LLM4FLw/o CodeNavigation        2 (-50.0%)        3 (-25.0%)        3 (-25.0%)        3 (-25.0%)
       LLM4FLw/o Divide&Conquer        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
       LLM4FLw/o Reflection        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
       LLM4FLw/o CoT        3 (-25.0%)        4 (0.0%)        4 (0.0%)        4 (0.0%)
       Jsoup (93)        LLM4FL        41        56        60        60
       LLM4FLw/o CodeNavigation        40 (-2.44%)        55 (-1.79%)        58 (-3.33%)        58 (-3.33%)
       LLM4FLw/o Divide&Conquer        36 (-12.2%)        50 (-10.71%)        51 (-15.0%)        51 (-15.0%)
       LLM4FLw/o Reflection        38 (-7.32%)        53 (-5.36%)        54 (-10.0%)        54 (-10.0%)
       LLM4FLw/o CoT        40 (-2.44%)        55 (-1.79%)        59 (-1.67%)        60 (0.0%)
       Lang (64)        LLM4FL        48        55        58        58
       LLM4FLw/o CodeNavigation        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o Divide&Conquer        43 (-10.42%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o Reflection        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o CoT        48 (0.0%)        55 (0.0%)        59 (1.72%)        59 (1.72%)
       Math (106)        LLM4FL        68        87        92        94
       LLM4FLw/o CodeNavigation        59 (-13.24%)        79 (-9.2%)        85 (-7.61%)        85 (-9.57%)
       LLM4FLw/o Divide&Conquer        49 (-27.94%)        68 (-21.84%)        70 (-23.91%)        78 (-17.02%)
       LLM4FLw/o Reflection        57 (-16.18%)        80 (-8.05%)        85 (-7.61%)        85 (-9.57%)
       LLM4FLw/o CoT        68 (0.0%)        87 (0.0%)        91 (-1.09%)        94 (0.0%)
       Mockito (38)        LLM4FL        21        22        26        27
       LLM4FLw/o CodeNavigation        18 (-14.29%)        23 (4.55%)        23 (-11.54%)        23 (-14.81%)
       LLM4FLw/o Divide&Conquer        15 (-28.57%)        20 (-9.09%)        21 (-19.23%)        21 (-22.22%)
       LLM4FLw/o Reflection        18 (-14.29%)        20 (-9.09%)        20 (-23.08%)        22 (-18.52%)
       LLM4FLw/o CoT        21 (0.0%)        22 (0.0%)        26 (0.0%)        27 (0.0%)
       Time (26)        LLM4FL        14        21        22        23
       LLM4FLw/o CodeNavigation        12 (-14.29%)        20 (-4.76%)        22 (0.0%)        22 (-4.35%)
       LLM4FLw/o Divide&Conquer        10 (-28.57%)        16 (-23.81%)        17 (-22.73%)        20 (-13.04%)
       LLM4FLw/o Reflection        12 (-14.29%)        19 (-9.52%)        19 (-13.64%)        22 (-4.35%)
       LLM4FLw/o CoT        14 (0.0%)        22 (4.76%)        23 (4.55%)        23 (0.0%)
       Total (675)        LLM4FL        327        425        473        494
       LLM4FLw/o CodeNavigation        273 (-16.51%)        378 (-11.06%)        409 (-13.53%)        409 (-17.21%)
       LLM4FLw/o Divide&Conquer        251 (-23.24%)        341 (-19.76%)        365 (-22.83%)        381 (-22.87%)
       LLM4FLw/o Reflection        290 (-11.31%)        400 (-5.88%)        436 (-7.82%)        459 (-7.09%)
       LLM4FLw/o CoT        325 (-0.61%)        424 (-0.24%)        473 (0.0%)        495 (0.2%)
Results. While all components help improve the results, including coverage grouping and prompt chaining provide the largest improvement to fault localization results (23% and 17% in Top-1). Table 4 shows the Top-1, 3, 5, and 10 scores when each component is removed. Removing coverage splitting has the largest overall impact across all scores, reducing Top-N by 19% to 23%. Removing prompt chaining has the second largest impact (11% to 17%). At the individual project level, these two components also have the largest impact in Top-1 in 9/13 studied projects. Our finding shows that employing sorted coverage grouping following the divide and conquer technique and agent communication significantly improves fault localization results. Future research should consider these techniques when designing fault localization techniques.
Although there is no oracle during the fault localization process, asking LLMs to self-reflection still helps improve the overall Top-1 by 11%. LLMs often suffer from hallucinations, especially when there is a lack of feedback from external oracles (Xu et al., 2024; Huang et al., 2023). Even though we did not provide any groundtruth or external feedback to LLM, we find that self-reflection is still effective in improving fault localization results. Self-reflect brings 6% to 11% improvement across the Top-N metrics. Our finding suggests that future studies should consider self-reflection even if there is no external feedback. Interestingly, we did not see large differences after removing chain-of-thought reasoning via test generation. In some cases, removing it even slightly improves the fault localization results. One reason may be that LLM4FL already includes self-reflection to make LLM think deeply about the result, and adding a chain-of-thought did not further improve the result. Our finding highlights the effectiveness of self-reflection, which should be considered in future fault localization results.
The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage splitting and prompt chaining having the largest positive impact. Removing these components leads to significant declines in accuracy, confirming their crucial role in handling token limitations and enabling more effective multi-agent collaboration.
4.3. RQ3: Does order matter in the initial list of methods provided to the LLM?
Motivation. LLM4FL divides the coverage data into segments to address the token size limitation of LLMs. We sort the methods using the Ochiai scores before segmentation, though different sorting mechanisms may affect the final fault localization result. Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024) observes that the order of premises affects LLM’s results. However, whether this effect extends to software engineering tasks, particularly fault localization, remains unclear. Hence, in this research question, we investigate whether the order of methods within the segments affects the LLM’s fault localization performance.
Approach. To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph. Each strategy provides a different way to sort the methods before they are segmented for further analysis.
LLM4FLExecution: In this baseline approach, we use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012). This default list represents the natural execution order of the methods, with no explicit ranking or prioritization. By providing the LLM with methods based on the execution sequence, we establish a control case to measure its performance without any ranking influence. The result shows how the order of premises in reasoning tasks can lead to varied LLM performance in software engineering context (Chen et al., 2024).
LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods during the segmentation process. Ochiai is unsupervised and is efficient to compute. We hypothesize that providing the LLM with methods sorted by their suspiciousness score will lead to more effective fault localization, as the model will focus on the most likely faulty candidates earlier in the process.
LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods during the segmentation process. DepGraph ranks methods based on structural code dependencies and code change history. As shown in RQ1, DepGraph shows the highest fault localization accuracy among all the techniques, surpassing LLM4FLOchiai. By examining the fault localization result after sorting the methods using DepGraph’s scores, we can better study if the initial order affects LLM’s results, even though LLM eventually visits all the methods.
@@ 4.3. RQ3: Does order matter in the initial list of methods provided to the LLM?
Results. Method ordering has a significant impact on LLM’s fault localization result, with up to 22% difference in Top-1 (from 299 to 366). Figure 4 shows the fault localization results when using different sorting strategies. When methods were presented in an execution-based order, LLM4FLExecution achieved a Top-1 score of 299, gradually increasing to 402 for Top-3, 431 for Top-5, and reaching 462 for Top-10. This performance establishes a baseline, showing how the LLM behaves without strategic ordering. However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements. LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution. The improvements also extended to other metrics, with the Top-3 score increasing to 421, the Top-5 reaching 469, and the Top-10 rising to 490.
LLM4FLDepGraph provides further improvement to the already-promising result of DepGraph, indicating method ordering is critical to LLM4FL, or LLM-based fault localization in general. LLM4FLDepGraph achieved the highest Top-1 score of 366, which is significantly outperforming both LLM4FLExecution and LLM4FLOchiai. The improvement was consistent across all the metrics. For Top-3, the score rose to 492, providing a substantial boost over LLM4FLOchiai. Similarly, LLM4FLDepGraph excelled in the Top-5 and Top-10 categories, reaching 543 and 548, respectively. We also see that LLM4FLDepGraph has better Top-1, 3, and 5 scores compared to DepGraph. This consistent improvement underscores the importance of method ordering in enhancing the accuracy of LLM-based fault localization. This underscores the importance of method ordering for LLM-based fault localization, showing that lightweight techniques like Ochiai can achieve substantial performance gains without the heavy computational burden. Striking this balance between accuracy and efficiency makes these approaches particularly suitable for a wide range of fault localization tasks. However, it is important to recognize that LLM4FLDepGraph is computationally expensive due to its reliance on graph neural networks (GNNs). While LLM4FLDepGraph delivers top-tier results, LLM4FLOchiai also offers a strong, efficient alternative, delivering good localization accuracy at a fraction of the computational cost due to the unsupervised nature of Ochiai. In other words, LLM4FLOchiai may be more easily adapted in practice.
Nevertheless, our finding establishes a new research direction for LLM-based fault localization. It demonstrates that intelligent method ordering strategies significantly impact the result of LLM-based fault localization. This approach opens up further opportunities for optimizing LLM-based fault localization by exploring more advanced ordering techniques and how different premises of ordering affect other software engineering tasks.
The findings highlight that method ordering plays a crucial role in improving the performance of LLM-based fault localization, with a difference of up to 22% in Top-1 scores. While LLM4FLDepGraph delivers the best results, the lightweight LLM4FLOchiai offers a highly efficient alternative, providing significant accuracy gains with far lower computational costs, making it more practical for real-world adoption.
@@ 5. Threats To Validity
Internal Validity. A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to data similar to the benchmark or specific fault localization cases during training. This could result in artificially inflated performance as the model may have prior knowledge of the evaluation data. Nevertheless, we mitigate this risk by preventing any content from being entered into ChatGPT that is related to the project name, the human-written bug report, or the bug ID.
External Validity. The primary threat to external validity is the generalizability of our results. Our evaluation is based on Defects4J, a well-established dataset in the software engineering community. Although this dataset includes real-world bugs, the systems studied are primarily Java-based, and it remains uncertain whether our findings will generalize to other programming languages or domains.
Construct Validity. Construct validity relates to whether the metrics we used accurately measure the performance of fault localization techniques. We used widely accepted Top-N metrics, which are commonly utilized in prior fault localization studies. However, our results are based on the assumption that developers primarily focus on the top-ranked faulty methods. Although this assumption aligns with previous research, different development practices could influence the effectiveness of our approach.
@@ 6. Conclusion
In this paper, we introduced LLM4FL, an LLM-agent-based fault localization technique that addresses the challenges of traditional fault localization methods and the limitations of current LLM-based approaches. By combining SBFL rankings with a divide-and-conquer strategy, LLM4FL groups large coverage data into manageable pieces, enabling effective analysis within the token limitations of LLMs. The use of multiple agents, prompt chaining, and self-reflection techniques allows for iterative refinement of fault localization, while chain-of-thought reasoning further enhances accuracy by generating potential fixes and re-rank the methods. Our evaluation on the Defects4J (V2.0.0) benchmark demonstrated that LLM4FL outperforms existing LLM-based techniques, achieving 19.27% higher Top-1 accuracy compared to AutoFL, and surpasses supervised approaches such as DeepFL and Grace without task-specific training. We also found that key components like coverage splitting and prompt chaining are essential to LLM4FL’s success, with method ordering playing a significant role in performance, improving Top-1 accuracy by up to 22%. Overall, LLM4FL presents a scalable and efficient solution for fault localization, providing improved accuracy and reducing computational costs, making it practical for real-world software projects. Future work will explore expanding LLM4FL’s capabilities for larger and more diverse codebases, further refining the agent collaboration and reasoning mechanisms.

### v2

@@ 4. STUDY DESIGN AND RESULTS
Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parentheses show the percentage changes compared to LLM4FL with all the components.
Techniques Top-1 Top-3 Top-5 Top-10
LLM4FL 327 425 473 494
LLM4FLw/o CodeNav 273 (-16.51%) 378 (-11.06%) 409 (-13.53%) 409 (-17.21%)
LLM4FLw/o Division 251 (-23.24%) 341 (-19.76%) 365 (-22.83%) 381 (-22.87%)
LLM4FLw/o Reflexion 290 (-11.31%) 400 (-5.88%) 436 (-7.82%) 459 (-7.09%)
Results. While all components help improve the results, including coverage division and code navigation provide the largest improvement to fault localization results (23% and 17% in Top-1). Table 4 shows the Top-1, 3, 5, and 10 scores when each component is removed. Removing coverage divisions has the largest overall impact across all scores, reducing Top-N by 19% to 23%. Removing prompt chaining has the second largest impact (11% to 17%). At the individual project level, these two components also have the largest impact in Top-1 in 9/13 studied projects. Our finding shows that employing sorted coverage grouping following the divide and conquer technique and agent communication significantly improves fault localization results. Future research should consider these techniques when designing fault localization techniques.
Although there is no oracle during the fault localization process, asking LLMs to self-reflect still helps improve the overall Top-1 by 11%. LLMs often suffer from hallucinations, especially when there is a lack of feedback from external oracles (Xu et al., 2024; Huang et al., 2023). Even though we did not provide any groundtruth or external feedback to LLM, we found that reflexion is still effective in improving fault localization results. We speculate that verbal reinforcement learning helps the model improve its results by revisiting the suspicious methods it ranked earlier, creating a feedback loop. In this process, we analyze the results for each group of methods and combine these group-specific results into a final ranking. The model can then spot mistakes or gaps in its logic, leading to better results. Self-reflection brings 6% to 11% improvement across the Top-N metrics. Our finding suggests that future studies should consider self-reflection even if there is no external feedback. Our finding highlights the effectiveness of self-reflection, which should be considered in future fault localization results.
The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage division and code navigation having the largest positive impact. Removing these components leads to significant declines in accuracy, confirming their role in finding faults in a large code-base.
RQ3: Does order matter in the initial list of methods provided to the LLM?
Motivation. LLM4FL divides the coverage data into different divisions using an order-aware division strategy to address the token size limitation of LLMs. We sort the methods using the Ochiai scores before the division, though different sorting mechanisms may affect the final fault localization result. Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024a) observes that the order of premises affects LLM’s results. However, whether this effect extends to software engineering tasks, particularly fault localization, remains unclear. Hence, in this research question, we investigate whether the order of methods within the groups affects the LLM’s fault localization performance.
Approach. To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph to sort the methods before the divide-and-conquer step.
LLM4FLExecution: We use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012). This default list represents the natural execution order of the methods, with no explicit ranking or prioritization. By providing the LLM with methods based on the execution sequence, we establish a control case to measure its performance without any ranking influence.
LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods before the divide-and-conquer process. Ochiai is unsupervised and is efficient to compute. We hypothesize that providing the LLM with methods sorted by their suspiciousness score will lead to more effective fault localization, as the model will focus on the most likely faulty candidates earlier in the process.
LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods. DepGraph ranks methods based on structural code dependencies and code change history. As shown in RQ1, DepGraph shows the highest fault localization accuracy among all the techniques, surpassing LLM4FLOchiai. By examining the fault localization result after sorting the methods using DepGraph’s scores, we can better study if the initial order affects LLM’s results, even though LLM eventually visits all the methods.
@@
Results. Method ordering has a significant impact on LLM’s fault localization result, with up to 22% difference in Top-1 (from 299 to 366). Figure 3 shows the fault localization results when using different sorting strategies. When methods were presented in the execution order, LLM4FLExecution achieved a Top-1 score of 299, 402 for Top-3, 431 for Top-5, and 462 for Top-10. This performance establishes a baseline, showing how the LLM behaves without strategic ordering. However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements across all Top-N, where LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution.
LLM4FLDepGraph provides further improvement to the already-promising result of DepGraph, indicating method ordering is critical to LLM4FL, or LLM-based fault localization in general. LLM4FLDepGraph achieved the highest Top-1 score of 366, which significantly outperforms both LLM4FLExecution and LLM4FLOchiai. The improvement was consistent across all the metrics. We also see that LLM4FLDepGraph has better Top-1, 3, and 5 scores compared to DepGraph. This consistent improvement underscores the importance of method ordering in enhancing the accuracy of LLM-based fault localization. Namely, if the initial order is closer to the group truth, the final localization result also tends to be more accurate.
Our finding establishes a new research direction for LLM-based fault localization, or any software engineering tasks that take a list of software artifacts as input. Future studies may study how different premises of ordering affect other software engineering tasks, and how to combine traditional software engineering techniques to pre-process LLM’s input to further improve the results.
The initial method ordering significantly impacts the accuracy of LLM-based fault localization, with Top-1 scores varying by up to 22%. Future research should explore different ordering strategies and how traditional software engineering techniques can be integrated to optimize LLM performance further.
@@ 5. Threats to Validity
Internal Validity. A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to the benchmark data during training. Nevertheless, Ramos et al. (Ramos et al., 2024) found that newer and larger models trained on larger datasets exhibit limited evidence of leakage for defect benchmarks. Since we utilize GPT-4o-mini, a large model trained using a tremendous amount of data, it likely shares similar characteristics in reducing memorization risks. Additionally, following prior work, we ensure no content related to the project name, human-written bug report, or bug ID is entered into ChatGPT to minimize the risk of data memorization (Qin et al., 2024).
External Validity. The primary threat to external validity is the generalizability of our results. Our evaluation is based on Defects4J, a well-established dataset in the software engineering community. Although this dataset includes real-world bugs, the systems studied are primarily Java-based. Future studies may extend our study to other programming languages or domains.
Construct Validity. Construct validity relates to whether the metrics we used accurately measure the performance of fault localization techniques. We used widely accepted Top-N metrics commonly utilized in prior fault localization studies. However, our results are based on the assumption that developers primarily focus on the top-ranked faulty methods. Although this assumption aligns with previous research, different development practices could influence the effectiveness of our approach.
@@ 6. Conclusion
In this paper, we introduced LLM4FL, an LLM-agent-based fault localization approach. It utilizes multiple specialized LLM agents, including Context, Debugger, and Reviewer, to iteratively refine and improve the accuracy of fault localization through order-aware prioritization, graph-based code navigation, and self-reflection using verbal reinforcement learning. Evaluated on the Defects4J (V2.0.0) benchmark, LLM4FL demonstrated significant improvements over existing approaches, achieving an 18.55% increase in Top-1 accuracy compared to AutoFL and 4.82% improvement over SoapFL. Further enhancements, including coverage segmentation and iterative refinement, increased accuracy by up to 22%. Future work will explore expanding LLM4FL’s capabilities for larger and more diverse codebases, further refining the agent collaboration and reasoning mechanisms.